OpenStudy (jakelee):

One number is 25 more than a second number. The sum of these two numbers is 106. Find the numbers. p.s define the variable, and write and equation

OpenStudy (zzr0ck3r):

let \(x\) be our first number and \(y\) our second "One number is 25 more than a second number" \[x=y+25\] " The sum of these two numbers is 106." \[x+y=106\] Plug the first equation into the second equation for \(x\).

OpenStudy (madhu.mukherjee.946):

let the first no be x

OpenStudy (madhu.mukherjee.946):

then the second number is 25 +x

OpenStudy (madhu.mukherjee.946):

so,x+x+25=106 2x=81 x=81/2

OpenStudy (jakelee):

thanks to both of you for the help

OpenStudy (madhu.mukherjee.946):

i think you should not consider the other number as y coz u see the statement ,the other number is 25 more than the second number so,it should be x+25

OpenStudy (madhu.mukherjee.946):

now you decide@jakelee

OpenStudy (jakelee):

oh okay thanks

OpenStudy (madhu.mukherjee.946):

your welcome

OpenStudy (triciaal):

both correct use one step or use 2 steps

OpenStudy (zzr0ck3r):

If you use the standard method you would name both variables \(x=y+25\) and \(x+y=106\) Plug in for \(x\) \((y+25)+y=106\) You will get the same answer...